### v1

@@ 1 Introduction
Recent years have witnessed the emergence of a combination of physics and data-driven models, offering a promising pathway to develop robust and interpretable CF models. These models strive to amalgamate domain-specific knowledge with deep learning prowess, thereby harnessing the strengths of both to create models adept at handling complex driving scenarios while adhering to real-world physics principles [25, 18, 19, 28]. However, these models’ significant limitation is their inability to comprehend a rational driver’s actions in full. Both learning-based models and PINN works have overlooked Rational Driving Constraints (RDCs) [29], a set of conditions on car following behavior that are generally accepted as necessary for sensible or rational driving behavior, and traditional physics-based models such as OVRV can only represent a simplified version of reality, excluding many complexities in driving behavior [15, 18, 19, 17].
@@
The remainder of this article unfolds as follows: In Section 2, we introduce relevant literature at the intersection of car-following modeling and deep learning. We then explore various types of car-following models and elaborate on our proposed methodology and model in Section 3. Section 4 presents the experimental data utilized in this study followed by the experiments in Section 5. Subsequently, Section 6 provides a comprehensive discussion of our numerical experiment results and their implications. Finally, we conclude the article in Section 7, where we discuss the study’s limitations and propose potential avenues for future research.
@@ 2.1 The Modeling of Driving Behavior
In response to these limitations, contemporary research is gravitating towards the integration of machine learning techniques for modeling driving behavior. These machine learning models, capable of learning directly from data, hold the potential to adapt to the patterns and variability inherent in individual drivers’ behavior. This shift has spurred the creation of innovative models such as the Physics-Informed Neural Networks (PINN) based car-following models
@@ 3 Modeling Car-following Behavior
In this section, we propose a rational neural network-based car-following model designed to meet the RDCs while being capable of adapting to varying driving conditions and unique driver behaviors. We will detail the model’s architecture, data processing, training process, and alignment with the RDC. Before delving into the proposed model, we first present an overview of the car-following model and three distinct types of modeling techniques. The classification of the driving policy mapping models includes physics-based, deep learning (DL)-based, and physics-guided AI car-following models.
@@
3.1 Physics-based Car-following Model
The input of the Physics-based Car-following Model (Phy-CFM) includes all possible signals $s(t)$ at time $t$ from neighboring vehicles in the traffic environment, and its output is the acceleration $a(t+\Delta t)$ at the subsequent time step. The Phy-CFM model aims to find an optimal set of parameters $\theta^{*}$ that best fits real-world data to minimize the error between a simulated trajectory for a following vehicle and the observed vehicle trajectory as measured in a dataset. One approach that has proven to be successful at finding best-fit model parameter values is to minimize the mean square error (MSE) in inter-vehicle spacing between the simulated car following trajectory and the inter-vehicle spacing in the experimental data [43, 20]. The procedure can be written as:
@@ 3.2 Deep Learning based Car-following Model
The application of machine learning techniques in modeling, classifying, and simulating car-following behaviors has garnered significant attention in recent years [15, 44, 45, 20]. These advanced systems have demonstrated superior performance over traditional physical models. However, they often grapple with challenges of interpretability and may exhibit limitations in handling scenarios that diverge from their training datasets.
@@ 3.3 Physics-guided AI Car-following Model
A major shortcoming of learning-based models is their inability to consistently yield physically plausible outputs. This issue could lead to unsafe and irrational driving behaviors resulting in vehicle collisions in microscopic traffic simulations [46]. However, the recent development of physics-informed neural networks (PINNs) offers a promising path toward a data-driven modeling paradigm that leverages the strengths of both physical and deep learning models [25]. A step in this direction is the car-following model proposed by Mo et al., which combines physical models with learning-based techniques [18, 19]. The key distinction between PINN-CFM and purely neural network-based CFM lies in the loss function, modified to include a weighted sum of two components: the data discrepancy loss $MSE_{NN}$ calculated by the neural network, and the physics discrepancy estimated using the physical car-following model $MSE_{Phy}$.
@@
$=\alpha\cdot\left(\frac{1}{N}\sum_{i=1}^{N}\left\|\hat{a}{(i)}-{a}_{NN}^{(i)}\right\|^{2}\right)+(1-\alpha)\cdot\left(\frac{1}{N}\sum_{i=1}^{N}\left\|{a}_{NN}^{(i)}-{a}_{Phy}^{(i)}\right\|^{2}\right)$
Here, $L_{\theta}$ denotes the loss function, $MSE_{NN}$ and $MSE_{Phy}$ represent the mean square errors for data discrepancy and physics discrepancy, respectively, $\alpha$ is the trade-off coefficient, and $N$ represents the size of observed data. The actual acceleration and its estimate for the observed state $i$ are denoted as $a{(i)}$. Finally, ${a}_{NN}^{(i)}$ and ${a}_{Phy}^{(i)}$ express the estimated accelerations from the neural network model and physics-based model, respectively.
@@ 3.4 Deep Learning-Based Car-Following Model Integration of Rational Driving Constraints
  • Decelerate when the speed is high. This condition is mathematically represented as a non-positive derivative of acceleration corresponding to speed, as shown in Equation 9.
@@
  • Accelerate when the spacing is large. The acceleration should increase as the spacing (distance from the preceding vehicle) increases. Mathematically, this is expressed as a non-negative derivative of acceleration corresponding to spacing:
@@
  • Accelerate when the relative speed is high. The acceleration should increase as the relative speed (speed difference with the preceding vehicle) increases. Mathematically, this is expressed as a non-negative derivative of acceleration corresponding to relative speed:
@@
Implementing these constraints in learning-based car-following models is essential as they warrant that the model’s predictions align with basic safe and rational driving principles. The proposed solution in the provided code introduces a novel approach to enforce RDCs by incorporating them into the loss function of the model. This effectively guides the model during training to learn predictions that satisfy RDCs.
@@
$=MSE(a_{pred},a_{true})+\lambda_{1}\cdot RDC_{speed}+\lambda_{2}\cdot RDC_{spacing}+\lambda_{3}\cdot RDC_{\text{relative speed}}$
@@
3:$Z_{\text{seq}}\leftarrow NN_{\text{seq}}(X_{\text{seq}})$
4:$Z_{\text{seq}}^{\prime}\leftarrow\text{process}(Z_{\text{seq}})$
5:$Z_{\text{phy}}^{\prime}\leftarrow\text{process}(X_{\text{phy}})$
6:$a_{\text{pred}}\leftarrow\text{combine}(Z_{\text{seq}}^{\prime},Z_{\text{phy}}^{\prime})$
@@
8:$RDC_{\text{speed}}\leftarrow\text{ReLU}\left(\frac{\partial a_{\text{pred}}}{\partial v}\right)$
9:$RDC_{\text{spacing}}\leftarrow\text{ReLU}\left(-\frac{\partial a_{\text{pred}}}{\partial s}\right)$
10:$RDC_{\text{relative speed}}\leftarrow\text{ReLU}\left(-\frac{\partial a_{\text{pred}}}{\partial r}\right)$
11:$\mathcal{L}\leftarrow MSE(a_{\text{pred}},a_{\text{true}})+\lambda\cdot(RDC_{\text{speed}}+RDC_{\text{spacing}}+RDC_{\text{relative speed}})$
@@
The details of the algorithm are presented in Algorithm 1. Here, $NN_{\text{seq}}$ represents the sequence handling neural network, $\text{process}(\cdot)$ signifies additional processing steps (that could vary depending on the exact network architecture used), and $\text{combine}(\cdot)$ refers to the operation that combines the processed sequence and physical inputs. The innovative use of two different inputs, $X_{\text{seq}}$ and $X_{\text{phy}}$, coupled with the application of the RDCs within the loss function, helps align the predictions of our model with the principles of vehicular movement, generating more realistic results. This integration of domain-specific knowledge and sequence data modeling is a crucial step toward the development of more reliable car-following models.
@@ 4 Data Description
Our analysis employs a principal dataset derived from a sequence of car-following experiments conducted by Gunter et al. [11]. This dataset is amassed using a variety of commercially available vehicles equipped with ACC systems. Each of these ACC-activated vehicles adheres to a uniform testing procedure, where a leading vehicle traverses at a pre-established speed sequence for a set duration at each pace. The ACC vehicle, while trailing the lead vehicle, has its ACC active throughout the experiment.
To ensure accurate position and speed data, each vehicle is fitted with high-precision GPS receivers. This allows for the calculation of the inter-vehicle gap and relative speed. Figure 1 provides a snapshot for one of the following vehicles under two conditions: minimum gap ACC setting, allowing the vehicle to maintain a close spacing to the one ahead, and maximum gap ACC setting, permitting greater spacing from the lead vehicle. This sample offers an illustrative example of leader-follower speed and spacing trajectories. The data collection operates at a 10 Hz sampling rate. Detailed information about the assorted set of vehicles utilized in Gunter et al.’s experiment [11] is provided in Table 1.
@@ 5.1 Physical Model
In this study, we employ the Optimal Velocity Relative Velocity (OVRV), an extended model to the FVD Model [27]. We select the OVRV model to depict the ACC vehicle driving behavior because it has demonstrated its ability to effectively encapsulate car-following dynamics of commercially-available ACC vehicles with a straightforward model [49, 12, 43].
@@ 5.3 PINN CF Model
$Loss_{PINN}=\alpha*MSE_{NN}+(1+\alpha)*MSE_{Physics}$ (26)
@@ 5.4 RACER Model
$=MSE(a_{pred},a_{true})+\lambda_{1}\cdot RDC_{speed}+\lambda_{2}\cdot RDC_{spacing}+\lambda_{3}\cdot RDC_{\text{relative speed}}$
@@ 6.2 Model Evaluation: ACC Vehicle with Minimum gap Setting
The performance of these models extends to predicting speed and spacing, which are crucial parameters that depend on the leading vehicle’s behavior. The OVRV model exhibits the weakest performance; its error margins for speed are at least twice as high as those of the RACER model and deteriorate further for spacing. Conversely, the PINN and NN models predict vehicle spacing with greater precision, as shown in Figure 4, and adapt more effectively to changes in the leading vehicle’s speed, shown in Figure 3. Nonetheless, the PINN and NN models’ performance in speed and spacing prediction is not as consistent as the RACER model, with RMSE values exceeding those of the RACER model, indicating room for improvement.
@@
We then scrutinize the predictions made by various deep learning models against testing data, specifically focusing on compliance with the RDCs. These constraints evaluate the model-predicted accelerations to three driving parameters: velocity ($v$), spacing ($s$), and relative speed ($r$).
Figure 5 presents the results for a baseline NN model. Intriguingly, the NN managed to fulfill the RDCs relative to speed and relative speed constraints but fell short in the context of spacing, exhibiting violations approximately 60% of the time, respectively. Furthermore, the model’s performance in relation to velocity did not improve significantly following a minor non-violation segment.
A different perspective is provided in Figure 6, which shows the performance of a PINN. Although the PINN model also complied fully with the RDCs of relative speed constraint, it exhibited a stark contrast when it came to velocity and spacing constraints, with an approximately 50% violation rate for both these parameters.
Finally, we turn our attention to our proposed model, the Rational Neural Network, the results of which are delineated in Figure 7. In stark contrast to the other models, this network demonstrates a flawless performance, recording zero RDC violations in terms of relative speed, velocity, or spacing. This outcome highlights the efficacy of our model in adhering to real-world driving principles, further reinforcing the benefits of integrating domain-specific knowledge into deep learning models for more realistic and reliable predictions.

### v2

@@ 1 Introduction
Recent years have witnessed the emergence of a combination of physics and data-driven models, offering a promising pathway to develop robust and interpretable CF models. These models strive to amalgamate domain-specific knowledge with deep learning prowess, thereby harnessing the strengths of both to create models adept at handling complex driving scenarios while adhering to real-world physics principles [25, 18, 19, 28, 29]. However, these models’ significant limitation is their inability to comprehend a rational driver’s actions in full. Both learning-based models and PINN works have overlooked Rational Driving Constraints (RDCs) [30], a set of conditions on car following behavior that are generally accepted as necessary for sensible or rational driving behavior, and traditional physics-based models such as OVRV can only represent a simplified version of reality, excluding many complexities in driving behavior [15, 18, 19, 17].
@@
The remainder of this article unfolds as follows: In Section 2, we introduce relevant literature at the intersection of car-following modeling and deep learning. We then explore various types of car-following models and elaborate on our proposed methodology and model in Section 3. Section 4 presents the experimental data utilized in this study followed by the experiments in Section 5. Subsequently, Section 6 provides a comprehensive discussion of our numerical experiment results and their implications. Finally, we conclude the article in Section 7, where we discuss the study’s limitations and propose potential avenues for future research. To provide a clear understanding of the research problem and methodology, we present a comprehensive list of notations used throughout this paper in Table 1.
Table 1: Nomenclature and acronyms used throughout the paper.
Symbol Description Unit
Variables and Parameters
$a_{t},\;\hat{a}_{t}$ True / predicted longitudinal acceleration m s-2
$a_{\text{NN}},\,a_{\text{Phy}}$ Acceleration predicted by NN / by physical model m s-2
$f_{\theta}(\cdot)$ Car-following mapping parameterized by $\theta$
$k_{1},k_{2}$ OVRV gain parameters
$N$ Number of data points in a batch
$s_{t}$ Space gap between follower and leader m
$\hat{s}(t)$ Measured inter-vehicle spacing m
$v_{t}$ Speed of follower vehicle m s-1
$\hat{v}_{\ell}(t)$, $\hat{v}_{f}(t)$ Measured lead / following vehicle speed m s-1
$X_{\text{seq}},X_{\text{phy}}$ Sequence input data / physical state input data
$\lambda_{1,2,3}$ Weights on RDC penalties in loss function
$\tau,\eta$ OVRV time constant / jam distance parameters s, m
$\theta$ CF model parameters
$\mathcal{O},\;\mathcal{A}$ Observation space / action space
Acronyms
ACC Adaptive Cruise Control
AV Automated Vehicle
CF Car-Following
LSTM Long Short-Term Memory network
MSE Mean Squared Error
OVRV Optimal Velocity–Relative Velocity model
PINN Physics-Informed Neural Network
RDC Rational Driving Constraint
RMSE Root Mean Squared Error
@@ 2.1 The Modeling of Driving Behavior
In response to these limitations, contemporary research is gravitating towards the integration of machine learning techniques for modeling driving behavior. These machine learning models, capable of learning directly from data, hold the potential to adapt to the patterns and variability inherent in individual drivers’ behavior. This shift has spurred the creation of innovative models such as the PINN based car-following models.
@@ 3 Modeling Car-following Behavior
In this section, we propose a rational neural network-based car-following model designed to meet the RDCs while being capable of adapting to varying driving conditions and unique driver behaviors. We will detail the model’s architecture, data processing, training process, and alignment with the RDC. Before delving into the proposed model, we first present an overview of the car-following model and three distinct types of modeling techniques. The classification of the driving policy mapping models includes engineering, deep learning (DL)-based, and physics-guided AI car-following models.
@@
3.1 Engineering Car-following Model
The input of the Physics-based Car-following Model (Phy-CFM) or Engineering Car-following Model includes all possible signals $s(t)$ at time $t$ from neighboring vehicles in the traffic environment, and its output is the acceleration $a(t+\Delta t)$ at the subsequent time step. The Phy-CFM model aims to find an optimal set of parameters $\theta^{*}$ that best fits real-world data to minimize the error between a simulated trajectory for a following vehicle and the observed vehicle trajectory as measured in a dataset. One approach that has proven to be successful at finding best-fit model parameter values is to minimize the mean square error (MSE) in inter-vehicle spacing between the simulated car following trajectory and the inter-vehicle spacing in the experimental data [44, 20]. The procedure can be written as:
@@ 3.2 Deep Learning based Car-following Model
The application of machine learning techniques in modeling, classifying, and simulating car-following behaviors has garnered significant attention in recent years [15, 45, 46, 20, 47]. These advanced systems have demonstrated superior performance over traditional physical models. However, they often grapple with challenges of interpretability and may exhibit limitations in handling scenarios that diverge from their training datasets.
@@ 3.3 Physics-guided AI Car-following Model
A major shortcoming of learning-based models is their inability to consistently yield physically plausible outputs. This issue could lead to unsafe and irrational driving behaviors resulting in vehicle collisions in microscopic traffic simulations [48]. However, the recent development of PINNs offers a promising path toward a data-driven modeling paradigm that leverages the strengths of both physical and deep learning models [25]. A step in this direction is the car-following model proposed by Mo et al., which combines physical models with learning-based techniques [18, 19]. The key distinction between PINN-CFM and purely neural network-based CFM lies in the loss function, modified to include a weighted sum of two components: the data discrepancy loss $MSE_{NN}$ calculated by the neural network, and the physics discrepancy estimated using the physical car-following model $MSE_{Phy}$.
@@
$=\alpha\cdot\left(\frac{1}{N}\sum_{i=1}^{N}\left\|\hat{a}{(i)}-{a}_{NN}^{(i)}\right\|^{2}\right)$
$+(1-\alpha)\cdot\left(\frac{1}{N}\sum_{i=1}^{N}\left\|{a}_{NN}^{(i)}-{a}_{Phy}^{(i)}\right\|^{2}\right)$
Here, $L_{\theta}$ denotes the loss function, $MSE_{NN}$ and $MSE_{Phy}$ represent the mean square errors for data discrepancy and physics discrepancy, respectively, $\alpha$ is the trade-off coefficient, and $N$ represents the size of observed data. The actual acceleration and its estimate for the observed state $i$ are denoted as $a{(i)}$. Finally, ${a}_{NN}^{(i)}$ and ${a}_{Phy}^{(i)}$ express the estimated accelerations from the neural network model and engineering models, respectively.
@@ 3.4 Deep Learning-Based Car-Following Model Integration of Rational Driving Constraints
  • Speed Constraint: Ensure that the derivative of acceleration with respect to speed is non-positive, indicating that the vehicle does not accelerate at higher speeds. This is mathematically represented as:
@@
  • Spacing Constraint: Ensure that the derivative of acceleration with respect to spacing is non-negative, indicating that the vehicle does not decelerate when the spacing increases. This is mathematically represented as:
@@
  • Relative Speed Constraint: Ensure that the derivative of acceleration with respect to relative speed is non-negative, indicating that the vehicle does not decelerate when the relative speed increases. This is mathematically represented as:
@@
Implementing these constraints in learning-based car-following models is essential as they warrant that the model’s predictions align with basic safe and rational driving principles. The proposed solution in the provided code introduces a novel approach to enforce RDCs by incorporating them into the loss function of the model. This effectively guides the model during training in learning predictions that satisfy RDCs.
@@
$=MSE(a_{pred},a_{true})+\lambda_{1}\cdot RDC_{speed}$
$+\lambda_{2}\cdot RDC_{spacing}+\lambda_{3}\cdot RDC_{\text{relative speed}}$
@@
The overall architecture of the RACER model is illustrated in Fig. 1. As shown, the model processes two distinct data inputs: sequential data ($X_{\text{seq}}$) and physical state data ($X_{\text{phy}}$). The process begins with (1) input encoding of this data, which then flows into (2) the neural network processing core. This pathway generates (3) an acceleration prediction ($a_{\text{pred}}$). Simultaneously, this prediction, along with current state variables, is used for (4) derivative computation, enabling the calculation of both the MSE loss and the RDC loss. These are combined into (5) a composite loss, which is then used for (6) backpropagation to update the model parameters. This end-to-end process, guided by both forward data flow (solid arrows) and gradient/derivative information (dashed arrows), ensures both predictive accuracy and rational driving constraint compliance.
@@
3:$Z_{\text{seq}}\leftarrow NN_{\text{seq}}(X_{\text{seq}})$triangleright LSTM encoding
4:$Z_{\text{seq}}^{\prime}\leftarrow\text{process}(Z_{\text{seq}})$triangleright Further processing
5:$Z_{\text{phy}}^{\prime}\leftarrow\text{process}(X_{\text{phy}})$triangleright Physical state processing
6:$a_{\text{pred}}\leftarrow\text{combine}(Z_{\text{seq}}^{\prime},Z_{\text{phy}}^{\prime})$triangleright Combine representations
@@
8:$RDC_{\text{speed}}\leftarrow\text{ReLU}\left(\frac{\partial a_{\text{pred}}}{\partial v}\right)$triangleright Speed constraint
9:$RDC_{\text{spacing}}\leftarrow\text{ReLU}\left(-\frac{\partial a_{\text{pred}}}{\partial s}\right)$triangleright Spacing constraint
10:$RDC_{\text{rel\_speed}}\leftarrow\text{ReLU}\left(-\frac{\partial a_{\text{pred}}}{\partial r}\right)$triangleright Relative speed constraint
11:$\mathcal{L}\leftarrow\text{MSE}(a_{\text{pred}},a_{\text{true}})+\lambda_{1}\cdot RDC_{\text{speed}}+\lambda_{2}\cdot RDC_{\text{spacing}}+\lambda_{3}\cdot RDC_{\text{rel\_speed}}$triangleright Combined loss
@@
The details of the algorithm are presented in Algorithm 1 and Fig. 1. The algorithm operates through several key stages: $NN_{\text{seq}}$ represents the sequence handling neural network (line 3), $\text{process}(\cdot)$ in lines 4-5 signifies additional processing steps that may vary depending on the exact network architecture used, and $\text{combine}(\cdot)$ in line 6 refers to the operation that combines the processed sequence and physical inputs. The core innovation lies in using two distinct inputs: $X_{\text{seq}}$ (temporal sequence data) and $X_{\text{phy}}$ (current kinematic state data typically used in physical models). Lines 7-10 implement the RDC enforcement by computing partial derivatives and applying ReLU functions to penalize constraint violations. The final loss function (line 11) combines the MSE loss for prediction accuracy with the RDC loss for constraint compliance. This integration of domain-specific knowledge and sequence data modeling represents a crucial step toward developing more reliable car-following models.
Figure 1: Architecture of the RACER model showing end-to-end dataflow: (1) input encoding of sequence data ($X_{\text{seq}}$) and kinematic states ($X_{\text{phy}}$); (2) neural network processing; (3) acceleration prediction ($a_{\text{pred}}$); (4) derivative computation for RDC calculation; (5) composite loss formation; (6) backpropagation. Solid arrows indicate forward flow; dashed arrows show gradient computations.
@@ 4 Data Description
Our analysis employs a principal dataset derived from a sequence of car-following experiments conducted by Gunter et al. [11]. This dataset is amassed using a variety of commercially available vehicles equipped with ACC systems. Each of these ACC-activated vehicles adheres to a uniform testing procedure, where a leading vehicle traverses at a pre-established speed sequence for a set duration at each pace. The ACC vehicle, while trailing the lead vehicle, has its ACC active throughout the experiment. To accurately emulate car-following behavior, the acceleration of the trailing vehicle is determined by evaluating its speed change over a time interval, corresponding to some time steps as the data is captured at a rate of 10 Hz (0.1 s intervals). The whole dataset includes about 10 minutes of vehicle driving trajectory, resulting in more than 6000 rows of data.
To ensure accurate position and speed data, each vehicle is fitted with high-precision GPS receivers. This allows for the calculation of the inter-vehicle gap and relative speed. Fig. 2 provides a snapshot for one of the following vehicles under two conditions: minimum gap ACC setting, allowing the vehicle to maintain a close spacing to the one ahead, and maximum gap ACC setting, permitting greater spacing from the lead vehicle. This sample offers an illustrative example of leader-follower speed and spacing trajectories. The data collection operates at a 10 Hz sampling rate. Detailed information about the assorted set of vehicles utilized in Gunter et al.’s experiment [11] is provided in Table 2. We used 80% of the data for training, 10% for validation, and the remaining 10% for testing.
@@ 5.1 Physical Model
In this study, we employ the OVRV, an extended model to the FVD Model [27]. We select the OVRV model to depict the ACC vehicle driving behavior because it has demonstrated its ability to effectively encapsulate car-following dynamics of commercially-available ACC vehicles with a straightforward model [51, 12, 44]. Additionally, we have reviewed other relevant literature [52, 53] and implemented widely used physical models such as IDM and Gipps for comparison with the OVRV model, using the same data and calibration methods. For the Gipps model, which relies on speed control instead of acceleration control, we estimated acceleration using headway, as outlined in the literature [52, 53], to ensure a fair comparison. The results show that the OVRV model performs best in both spacing and speed metrics, making it our chosen baseline physical model.
@@ 5.3 PINN CF Model
$Loss_{PINN}=\alpha\cdot MSE_{NN}+(1-\alpha)\cdot MSE_{Physics}$ (26)
@@ 5.4 RACER Model
$=MSE(a_{pred},a_{true})+\lambda_{1}\cdot RDC_{speed}$
$+\lambda_{2}\cdot RDC_{spacing}+\lambda_{3}\cdot RDC_{\text{relative speed}}$
@@ 6.2 Model Evaluation: ACC Vehicle with Minimum gap Setting
The performance of these models extends to predicting speed and spacing, which are crucial parameters that depend on the leading vehicle’s behavior. The OVRV model exhibits the weakest performance; its error margins for speed are at least twice as high as those of the RACER model and deteriorate further for spacing. Conversely, the PINN and NN models predict vehicle spacing with greater precision, as shown in Fig. 5, and adapt more effectively to changes in the leading vehicle’s speed, shown in Fig. 4. Nonetheless, the PINN and NN models’ performance in speed and spacing prediction is not as consistent as the RACER model, with RMSE values exceeding those of the RACER model, indicating room for improvement.
@@
We then scrutinize the predictions made by various deep learning models against testing data, specifically focusing on compliance with the RDCs. These constraints evaluate the model-predicted accelerations to three driving parameters: velocity ($v$), spacing ($s$), and relative speed ($r$). To accurately compute these constraints, we utilize PyTorch’s automatic differentiation function, which allows us to compute the precise gradients of acceleration with respect to each variable. This approach ensures that the derivatives are computed exactly at each time step, rather than approximating them numerically. As such, any violations in the RDCs are detected based on the multivariate derivatives across velocity, spacing, and relative speed.
Fig. 6 presents the results for a baseline NN model. Intriguingly, the NN managed to fulfill the RDCs relative to speed and relative speed constraints but fell short in the context of spacing, exhibiting violations approximately 60% of the time, respectively. These violations are highlighted in the figure using red lines to indicate where the partial derivative $\frac{da}{dv}\leq 0$, signaling a violation of the RDC. Additionally, the model’s performance with respect to velocity did not improve significantly, despite some minor non-violation segments.
A different perspective is provided in Fig. 7, which shows the performance of a PINN. Although the PINN model also complied fully with the RDCs of relative speed constraint, it exhibited a stark contrast when it came to velocity and spacing constraints, with an approximately 50% violation rate for both these parameters.
Finally, we turn our attention to our proposed model, the Rational Neural Network, the results of which are delineated in Fig. 8. Using the same automatic differentiation approach for detecting RDC violations, this network demonstrates flawless performance, with zero RDC violations for relative speed, velocity, or spacing. Unlike the other models, this network successfully adheres to the constraints across all variables. This outcome highlights the efficacy of our model in adhering to real-world driving principles, further reinforcing the benefits of integrating domain-specific knowledge into deep learning models for more realistic and reliable predictions. The superior performance of RACER can be attributed to three key mechanisms: (i) the explicit enforcement of RDCs prevents the model from learning physically implausible behaviors that lead to prediction errors, (ii) the dual-input architecture ($X_{\text{seq}}$ and $X_{\text{phy}}$) allows the model to leverage both temporal patterns and instantaneous physical constraints simultaneously, and (iii) the gradient-based RDC enforcement through automatic differentiation ensures that the learned mapping respects fundamental driving principles at every prediction step, leading to more stable and realistic long-term trajectory simulation.
The OVRV model’s poor performance in spacing prediction (RMSE of 1.47 m vs. RACER’s 0.261 m) stems from its simplified linear relationships that cannot capture the complex, nonlinear dynamics present in real ACC behavior. The standard Neural Network’s high violation rates ( 60% for spacing constraints) occur because it optimizes purely for prediction accuracy without physical constraints, leading to mathematically optimal but physically implausible solutions that violate basic driving principles. The PINN model’s moderate violations ( 50%) result from the soft enforcement of physics through the loss function—while it encourages physics compliance, it cannot guarantee it, allowing violations when data fitting objectives conflict with physical principles.